# Bifurcation sequences in the secular 3D planetary 3-Body problem: a geometric approach

### v2

@@ 4.2 Secular normal form $\mathcal{H}_{int}^{(1)}$
Figure 8: Upper panels: the contour plots $\mathcal{H}^{(1)}_{int}(X_{2},Y_{2};\sigma_{0})=\mathcal{E}$, with $Y_{3}=0$, $\dot{Y_{3}}\geq 0$ for decreasing values of $\sigma_{0}$ from top to bottom, left to right. Lower panels: intersections of and tangencies between the circles $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ (blue) and of the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ in the plane $(\sigma_{1},\sigma_{3})$, for different values of the energy $\mathcal{E}$ and $\sigma_{0}$ fixed as in the upper frame corresponding to each lower frame. The points of tangency corresponding to CP1-type fixed points of the secular normal form are calculated by Eq. (32). The curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ passing through the tangency points are hyperbola-like (compare with the corresponding curves in Fig. 6, which are ellipse-like).
@@
Figure 10: Zoom of the second plot of the second row of Fig. 8 (i.e. of the intersection between $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ in the plane $(\sigma_{1},\sigma_{3})$ for $\sigma_{0}=1.023\cdot 10^{-2}$). The blue and red points are, respectively, unstable and stable since they are given by an inner and outer tangency between $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$. On the left we show the inner tangency between $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{1})}}^{(\sigma_{2}=0)}$ in $P_{1}$, while on the right the outer tangency between $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{2})}}^{(\sigma_{2}=0)}$ in $P_{2}$ (in our example $\mathcal{E}^{(P_{1})}=-3.78701$ and $\mathcal{E}^{(P_{2})}=-3.78861$).
The sequence of bifurcations found in Fig. 8 can be summarized as follows: The top-left frame shows the starting phase portrait of the (near-)planar regime with the usual tangency points A and B associated with anti-aligned and aligned ACR orbits. Decreasing the value of $\sigma_{0}$ (second frame of Fig. 8, for $\sigma_{0}=1.023\cdot 10^{-2}$) gives rise to a similar saddle-node bifurcation as in Fig. 6, with two new fixed points of the Poincaré map corresponding to periodic orbits $P_{1}$ and $P_{2}$. However, due to the fact that the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ are now hyperbola-like, we note that the stability type of each of the orbits $P_{1}$ or $P_{2}$ is inverted with respect to Fig. 6. Thus, immediately after the bifurcation the tangent points belong to two nearby hyperbola-like curves, of which the curve containing the point $P_{1}$ yields an inner tangency (hence $P_{1}$ is unstable), while the curve containing the point $P_{2}$ yields an outer tangency (hence $P_{2}$ is unstable) (see a zoom in Fig. 10). Decreasing further the value of $\sigma_{0}$, we arrive at a critical value ($\sigma_{0}\simeq 1.012\cdot 10^{-2}$, see fourth panel in Fig. 8) where two independent hyperbola-like branches of the curve $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{2})}}^{(\sigma_{2}=0)}$ join each other into a saddle point $F$. This point is outside the disc limited by the curve $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$, hence it represents no possible new fixed point of the flows. Decreasing, however, $\sigma_{0}$ still further, the saddle point $F$ moves towards the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$, and crosses it at a critical value of $\sigma_{0}$ which lies is in the interval $8.94\cdot 10^{-3}<\sigma_{0}<9.91\cdot 10^{-3}$ (see the transition between the fifth and sixth panel of Fig. 8). At the crossing value we then have an inverse pitchfork bifurcation, whereby the point $P_{1}$ turns from unstable to stable, while for values of $\sigma_{0}$ below the critical one the saddle moves within the disc, giving rise to two unstable fixed points $F_{1}$ and $F_{2}$ of the CP2-type. For still smaller values of $\sigma_{0}$, the saddle point $F$ moves further inside the disc towards the left. Figure 11 gives the 3D picture of the intersection of the surfaces $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(F)}}$ with the sphere $\mathcal{S}_{\sigma_{0}}$ for $\sigma_{0}=8.94\cdot 10^{-3}$ and for $\sigma_{0}=8.14\cdot 10^{-3}$, corresponding to the sixth and seven panels in Fig. 8. Reducing $\sigma_{0}$ further, we then arrive at a second critical value lying in the interval $7.7\cdot 10^{-3}<\sigma_{0}<8.14\cdot 10^{-3}$ (see the transition between the seventh and eighth panel of Fig. 8), where the saddle point $F$ crosses for a second time the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$, passing now from its interior to its exterior. At the second crossing we then have a second inverse pitchfork bifurcation turning the point B from stable to unstable. From this point on, decreasing $\sigma_{0}$ brings the saddle point $F$ outside the disc, hence no longer yielding feasible fixed points of the system. In the same time, the ACR points B and A are given by an inner and outer tangency, hence they are unstable and stable respectively. Finally, reducing further $\sigma_{0}$ brings the separatrices of the fixed point B very close to the tangency point A. Hence, the area occupied by the island of stability around the fixed point A is necessarily reduced, while the areas of the domains of stability around the fixed points $P_{1}$, and $P_{2}$ become large. As a consequence, the points $P_{1}$ and $P_{2}$ and their neighborhoods give now the dominant structures in the phase space (see 9th and 10th panel in Fig. 8), while the influence of the ACR orbits A, B is reduced. Then, at a third critical value of $\sigma_{0}$ in the interval $7.33\cdot 10^{-3}<\sigma_{0}<7.49\cdot 10^{-3}$ the points A, B, as well as the hyperbolas which contain them, collide. Then, the points A, B disappear altogether through an inverse saddle-node bifurcation (see the transition from the 11th to the 12th (last) panel in Fig. 8). From that point on, the only stable points surviving in the end of the sequence of bifurcations are $P_{1}$ and $P_{2}$.
@@
Figure 11: Graphical representation of the critical points of second kind $F_{1}$ and $F_{2}$. Left: 3D representation of the sphere $\mathcal{S}_{\sigma_{0}}$ and the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}_{1}^{(F)}}$, in the case $\sigma_{0}=8.94\cdot 10^{-3}$, where $\mathcal{E}_{1}^{(F)}$ represents the energy of the fixed points $F_{1}$ and $F_{2}$. We can observe that the surface of the energy pierces the sphere in the $F$-modes. Centre: As before, but in the case $\sigma_{0}=8.14\cdot 10^{-3}\,$. For a value of the energy, namely $\mathcal{E}_{2}^{(F)}$, $\mathcal{C}_{\sigma_{0},\mathcal{E}_{2}^{(F)}}$ pierces the sphere $\mathcal{S}_{\sigma_{0}}$ in the points $F_{1}$ and $F_{2}$ of coordinates, respectively, $(\sigma_{1}^{(F)},\sigma_{2}^{(F)},\sigma_{3}^{(F)})$ and $(\sigma_{1}^{(F)},-\sigma_{2}^{(F)},\sigma_{3}^{(F)})$. Right: The projection of the 3D central plot (case $\sigma_{0}=8.14\cdot 10^{-3}\,$) in the planes $\sigma_{2}=0$.
4.3 Secular normal form: octupole approximation ~$\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$
@@
Of all multipolar truncations of the secular Hamiltonian, the first non-integrable one is the octupolar approximation, corresponding to terms up to degree $N_{P}=3$ in the semi-major axis ratio $a_{2}/a_{3}$. In this case, assuming an expansion up to order $N_{bk}=4$ in the eccentricities leads to the possibility of computing a secular normal form as in subsection 4.2, whose expression is an exact quadratic form in the Hopf variables as in Eq. (34). Explicit expressions for the normal form coefficients can be computed leading to:
~$\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}=A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}+D(\sigma_{0})\sigma_{1}+E(\sigma_{0})\sigma_{3}+F(\sigma_{0})\,,$ (52)
where the coefficients (apart from constants) are:
$A$ $=0\,,$
$C$ $=-\frac{2025a_{2}^{9/2}\mathrm{AMD}\left(\sqrt{a_{2}}m_{2}+\sqrt{a_{3}}m_{3}\right)^{2}}{256a_{3}^{19/2}{m_{2}}{m_{3}}{m_{0}}^{2}(a+b)}+\frac{225{a_{2}}^{2}\mathrm{AMD}}{64{a_{3}}^{6}{m_{0}}^{2}a}\left(\sqrt{\frac{a_{2}}{a_{3}}}\frac{m_{3}}{m_{2}}+\frac{{a_{2}}}{2{a_{3}}}+\frac{{m_{3}}^{2}}{2{m_{2}}^{2}}\right)$
@@
$B$ $=\frac{675{a_{2}}^{13/4}\mathrm{AMD}\left(\sqrt{{a_{2}}}{m_{2}}+\sqrt{{a_{3}}}{m_{3}}\right)^{2}}{256{a_{3}}^{33/4}{m_{0}}^{2}{m_{2}}^{3/2}\sqrt{{m_{3}}}(a+b)}\left(\frac{{b}}{{a}}+3\right)+\frac{5{a_{2}}^{9/4}}{128{a_{3}}^{17/4}{m_{0}}}\left(57\sqrt{\frac{{a_{2}m_{2}}}{a_{3}m_{3}}}-15\sqrt{\frac{{m_{3}}}{m_{2}}}\right)\,,$
$D_{1}$ $=\frac{2025{a_{2}}^{13/4}\mathrm{AMD}\left(\sqrt{{a_{2}}}{m_{2}}+\sqrt{{a_{3}}}{m_{3}}\right)^{2}}{256{a_{3}}^{33/4}{m_{0}}^{2}{m_{2}}^{3/2}\sqrt{{m_{3}}}(a+b)}\left(\frac{b}{a}+3\right)-\frac{105{a_{2}}^{9/4}}{128{a_{3}}^{17/4}{m_{0}}}\left(9\sqrt{\frac{{a_{2}m_{2}}}{a_{3}m_{3}}}+7\sqrt{\frac{{m_{3}}}{m_{2}}}\right)\,,$
$\Delta_{1}$ $=-\frac{675{a_{2}}^{13/4}\mathrm{AMD}^{2}\left(\sqrt{{a_{2}}}{m_{2}}+\sqrt{{a_{3}}}{m_{3}}\right)^{2}}{256{a_{3}}^{33/4}{m_{0}}^{2}{m_{2}}^{3/2}\sqrt{{m_{3}}}(a+b)}\left(\frac{b}{a}+3\right)+\frac{165{a_{2}}^{9/4}\mathrm{AMD}\left(\sqrt{{a_{2}}}{m_{2}}+\sqrt{{a_{3}}}{m_{3}}\right)}{32{a_{3}}^{19/4}{m_{0}}\sqrt{{m_{2}}}\sqrt{{m_{3}}}}$
@@
$D_{3}$ $=\frac{675{a_{2}}^{2}\mathrm{AMD}}{32{a_{3}}^{6}{m_{0}}^{2}a}\left(\sqrt{\frac{a_{2}}{a_{3}}}\frac{{m_{3}}}{{m_{2}}}+\frac{{a_{2}}}{2{a_{3}}}+\frac{{m_{3}}^{2}}{2{m_{2}}^{2}}\right)+\frac{3{a_{2}}}{4{a_{3}}^{3}{m_{0}}}\left(\frac{3{a_{2}}{m_{2}}}{{a_{3}}{m_{3}}}-\frac{7{m_{3}}}{4{m_{2}}}\right)\,,$
$\Delta_{3}$ $=\frac{3{a_{2}}^{3/2}\sqrt{\mathcal{G}}}{8{a_{3}}^{7/2}\sqrt{{m_{0}}}}\left(\sqrt{{a_{2}}}{m_{2}}-\sqrt{{a_{3}}}{m_{3}}\right)-\frac{225{a_{2}}^{2}\mathrm{AMD}^{2}}{32{a_{3}}^{6}{m_{0}}^{2}a}\left(\sqrt{\frac{a_{2}}{a_{3}}}\frac{{m_{3}}}{{m_{2}}}+\frac{{a_{2}}}{2a_{3}}+\frac{{m_{3}}^{2}}{2{m_{2}}^{2}}\right)$
$\phantom{=}+\frac{3{a_{2}}\mathrm{AMD}}{2{a_{3}}^{3}{m_{0}}}\left(\frac{m_{3}}{{m_{2}}}-\frac{{a_{2}}{m_{2}}}{{a_{3}}{m_{3}}}\right)\,.$
@@
In our numerical example, dropping constants, as well as an irrelevant term depending only on $\sigma_{0}$, the secular normal form in the octupole approximation resumes the form of Eq. (52) with
$A\!=0,$ $C\!=-0.0283636,$ $B\!=-0.029019,$
$D(\sigma_{0})\!=0.0011798-0.21118\,\sigma_{0},$ $E(\sigma_{0})\!=0.00153399-0.270077\,\sigma_{0}\,.$
@@
~$\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}=\mathcal{K}_{I}(\tilde{\sigma}_{1},\tilde{\sigma}_{3};\sigma_{0})=\mathcal{A}\tilde{\sigma}_{1}^{2}+\mathcal{C}\tilde{\sigma}_{3}^{2}+(\mathcal{D}_{1}{\sigma}_{0}+\Delta_{1})\tilde{\sigma}_{1}+(\mathcal{D}_{3}{\sigma}_{0}+\Delta_{3})\tilde{\sigma}_{3}$ (53)

### v3

@@ 4.2 Secular normal form $\mathcal{H}_{int}^{(1)}$
Figure 9: Upper panels: the contour plots $\mathcal{H}^{(1)}_{int}(X_{2},Y_{2};\sigma_{0})=\mathcal{E}$ (where $\mathcal{H}^{(1)}_{int}$ written in Eq. (59)), with $Y_{3}=0$, $\dot{Y_{3}}\geq 0$ for decreasing values of $\sigma_{0}$ from top to bottom, left to right. Lower panels: intersections of and tangencies between the circles $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ (blue) and of the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ in the plane $(\sigma_{1},\sigma_{3})$, for different values of the energy $\mathcal{E}$ and $\sigma_{0}$ fixed as in the upper frame corresponding to each lower frame. The points of tangency corresponding to CPI-type fixed points of the secular normal form are calculated by Eq. (36). The curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ passing through the tangency points are hyperbola-like (compare with the corresponding curves in Fig. 7, which are ellipse-like).
@@
Figure 11: Zoom of the second plot of the second row of Fig. 9 (i.e. of the intersection between $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ in the plane $(\sigma_{1},\sigma_{3})$ for $\sigma_{0}=1.023\cdot 10^{-2}$). The blue and red points are, respectively, unstable and stable since they are given by an inner and outer tangency between $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$. The left panel shows the inner tangency between $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{1})}}^{(\sigma_{2}=0)}$ in $P_{1}$. The middle panel shows the outer tangency between $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{2})}}^{(\sigma_{2}=0)}$ in $P_{2}$ (in our example $\mathcal{E}^{(P_{1})}=-3.78701$ and $\mathcal{E}^{(P_{2})}=-3.78861$). The right panel shows a zoom of the same plots in the domain around the points $P_{1}$ and $P_{2}$. Since for this value of $\sigma_{0}$ the saddle-node bifurcation has already occurred, the two hyperbolas are very close to each other but different, giving rise to two different tangencies at the points $P_{1}$ and $P_{2}$.
@@
Figure 12: Graphical representation of the critical points of second kind $F_{1}$ and $F_{2}$. Top left: 3D representation of the sphere $\mathcal{S}_{\sigma_{0}}$ and the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}_{1}^{(F)}}$, in the case $\sigma_{0}=8.94\cdot 10^{-3}$, where $\mathcal{E}_{1}^{(F)}$ represents the energy of the fixed points $F_{1}$ and $F_{2}$. We can observe that the surface of the energy pierces the sphere in the $F$-modes. Top right: As before, but in the case $\sigma_{0}=8.14\cdot 10^{-3}\,$. For a value of the energy, namely $\mathcal{E}_{2}^{(F)}$, $\mathcal{C}_{\sigma_{0},\mathcal{E}_{2}^{(F)}}$ pierces the sphere $\mathcal{S}_{\sigma_{0}}$ in the points $F_{1}$ and $F_{2}$ of coordinates, respectively, $(\sigma_{1}^{(F)},\sigma_{2}^{(F)},\sigma_{3}^{(F)})$ and $(\sigma_{1}^{(F)},-\sigma_{2}^{(F)},\sigma_{3}^{(F)})$. Bottom: The projection of the 3D top right plot (case $\sigma_{0}=8.14\cdot 10^{-3}\,$) in the planes $\sigma_{2}=0$.
The sequence of bifurcations found in Fig. 9 can be summarized as follows: The top-left frame shows the starting phase portrait of the (near-)planar regime with the usual tangency points A and B associated with anti-aligned and aligned ACR orbits. Decreasing the value of $\sigma_{0}$ (second frame of Fig. 9, for $\sigma_{0}=1.023\cdot 10^{-2}$) gives rise to a similar saddle-node bifurcation as in Fig. 7, with two new fixed points of the Poincaré map corresponding to periodic orbits $P_{1}$ and $P_{2}$. However, due to the fact that the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ are now hyperbola-like, we note that the stability type of each of the orbits $P_{1}$ or $P_{2}$ is inverted with respect to Fig. 7. Thus, immediately after the bifurcation the tangent points belong to two nearby hyperbola-like curves, of which the curve containing the point $P_{1}$ yields an inner tangency (hence $P_{1}$ is unstable), while the curve containing the point $P_{2}$ yields an outer tangency (hence $P_{2}$ is unstable) (see a zoom in Fig. 11). Decreasing further the value of $\sigma_{0}$, we arrive at a critical value ($\sigma_{0}\simeq 1.012\cdot 10^{-2}$, see fourth panel in Fig. 9) where two independent hyperbola-like branches of the curve $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{2})}}^{(\sigma_{2}=0)}$ join each other into a saddle point $F$. This point is outside the disc limited by the curve $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$, hence it represents no possible new fixed point of the flows. Decreasing, however, $\sigma_{0}$ still further, the saddle point $F$ moves towards the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$, and crosses it at a critical value of $\sigma_{0}$ which lies is in the interval $8.94\cdot 10^{-3}<\sigma_{0}<9.91\cdot 10^{-3}$ (see the transition between the fifth and sixth panel of Fig. 9). At the crossing value we then have an inverse pitchfork bifurcation, whereby the point $P_{1}$ turns from unstable to stable, while for values of $\sigma_{0}$ below the critical one the saddle moves within the disc, giving rise to two unstable fixed points $F_{1}$ and $F_{2}$ of the CPII-type. For still smaller values of $\sigma_{0}$, the saddle point $F$ moves further inside the disc towards the left. Figure 12 gives the 3D picture of the intersection of the surfaces $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(F)}}$ with the sphere $\mathcal{S}_{\sigma_{0}}$ for $\sigma_{0}=8.94\cdot 10^{-3}$ and for $\sigma_{0}=8.14\cdot 10^{-3}$, corresponding to the sixth and seven panels in Fig. 9. Reducing $\sigma_{0}$ further, we then arrive at a second critical value lying in the interval $7.7\cdot 10^{-3}<\sigma_{0}<8.14\cdot 10^{-3}$ (see the transition between the seventh and eighth panel of Fig. 9), where the saddle point $F$ crosses for a second time the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$, passing now from its interior to its exterior. At the second crossing we then have a second inverse pitchfork bifurcation turning the point B from stable to unstable. From this point on, decreasing $\sigma_{0}$ brings the saddle point $F$ outside the disc, hence no longer yielding feasible fixed points of the system. In the same time, the ACR points B and A are given by an inner and outer tangency, hence they are unstable and stable respectively. Finally, reducing further $\sigma_{0}$ brings the separatrices of the fixed point B very close to the tangency point A. Hence, the area occupied by the island of stability around the fixed point A is necessarily reduced, while the areas of the domains of stability around the fixed points $P_{1}$, and $P_{2}$ become large. As a consequence, the points $P_{1}$ and $P_{2}$ and their neighborhoods give now the dominant structures in the phase space (see 9th and 10th panel in Fig. 9), while the influence of the ACR orbits A, B is reduced. Then, at a third critical value of $\sigma_{0}$ in the interval $7.33\cdot 10^{-3}<\sigma_{0}<7.49\cdot 10^{-3}$ the points A, B, as well as the hyperbolas which contain them, collide. Then, the points A, B disappear altogether through an inverse saddle-node bifurcation (see the transition from the 11th to the 12th (last) panel in Fig. 9). From that point on, the only stable points surviving in the end of the sequence of bifurcations are $P_{1}$ and $P_{2}$.
As an overall comment, the sequence of bifurcations produced by the integrable Hamiltonian $\mathcal{H}_{int}$ differ considerably from those of the full secular Hamiltonian, the difference becoming more important as we approach towards the ‘Lidov-Kozai’ regime of dynamics. In particular, the full secular Hamiltonian $\mathcal{H}_{sec}$ contains a set of harmonics $\cos(k_{2}w_{2}+k_{3}w_{3})$ with $k_{2}+k_{3}\neq 0$, which are not present in the integrable approximation $\mathcal{H}_{int}$. Of the missing harmonics, the most important is $\cos(2w_{2})$ ($k_{2}=2$, $k_{3}=0$), which is the prevalent harmonic at the Lidov-Kozai regime.
4.3 Secular normal form: octupole approximation $\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$
@@
Of all multipolar truncations of the secular Hamiltonian, the first non-integrable one is the octupolar approximation, corresponding to terms up to degree $N_{P}=3$ in the semi-major axis ratio $a_{2}/a_{3}$. More precisely, starting from the octupole approximation of the Hamiltonian, namely $\mathcal{H}^{(N_{P}=3,N_{bk})}$ and decomposing it similarly as in Eq. (17), we can apply the normal form described in Section 4.2, and obtain an integrable approximation $\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk})}$ admitting $\sigma_{0}$ as a second integral. Moreover, in this case, assuming an expansion up to order $N_{bk}=4$ in the eccentricities leads to the possibility of computing a secular normal form as in subsection 4.2, whose expression is an exact quadratic form in the Hopf variables as in Eq. (38). In turn, this allows to apply the analytical formulæ reported in Section 3.4. Explicit expressions for the normal form coefficients can be computed, leading to a Hamiltonian of the same form of Eq. (38). More precisely:
$\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}={\tilde{A}}\sigma_{1}^{2}+{\tilde{C}}\sigma_{3}^{2}+{\tilde{B}}\sigma_{1}\sigma_{3}+{\tilde{D}}(\sigma_{0})\sigma_{1}+{\tilde{E}}(\sigma_{0})\sigma_{3}+{\tilde{F}}(\sigma_{0})\,$, (62)
where the coefficients (apart from additive constants) are:
${\tilde{A}}$ $=0\,$,
${\tilde{C}}$ $=-\frac{2025a_{2}^{9/2}\mathrm{AMD}\left(\sqrt{a_{2}}m_{2}+\sqrt{a_{3}}m_{3}\right)^{2}}{256a_{3}^{19/2}{m_{2}}{m_{3}}{m_{0}}^{2}(a+b)}+\frac{225{a_{2}}^{2}\mathrm{AMD}}{64{a_{3}}^{6}{m_{0}}^{2}a}\left(\sqrt{\frac{a_{2}}{a_{3}}}\frac{m_{3}}{m_{2}}+\frac{{a_{2}}}{2{a_{3}}}+\frac{{m_{3}}^{2}}{2{m_{2}}^{2}}\right)$
@@
${\tilde{B}}$ $=\frac{675{a_{2}}^{13/4}\mathrm{AMD}\left(\sqrt{{a_{2}}}{m_{2}}+\sqrt{{a_{3}}}{m_{3}}\right)^{2}}{256{a_{3}}^{33/4}{m_{0}}^{2}{m_{2}}^{3/2}\sqrt{{m_{3}}}(a+b)}\left(\frac{{b}}{{a}}+3\right)+\frac{5{a_{2}}^{9/4}}{128{a_{3}}^{17/4}{m_{0}}}\left(57\sqrt{\frac{{a_{2}m_{2}}}{a_{3}m_{3}}}-15\sqrt{\frac{{m_{3}}}{m_{2}}}\right)\,$,
${\tilde{D}}_{1}$ $=\frac{2025{a_{2}}^{13/4}\mathrm{AMD}\left(\sqrt{{a_{2}}}{m_{2}}+\sqrt{{a_{3}}}{m_{3}}\right)^{2}}{256{a_{3}}^{33/4}{m_{0}}^{2}{m_{2}}^{3/2}\sqrt{{m_{3}}}(a+b)}\left(\frac{b}{a}+3\right)-\frac{105{a_{2}}^{9/4}}{128{a_{3}}^{17/4}{m_{0}}}\left(9\sqrt{\frac{{a_{2}m_{2}}}{a_{3}m_{3}}}+7\sqrt{\frac{{m_{3}}}{m_{2}}}\right)\,$,
${\tilde{\Delta}}_{1}$ $=-\frac{675{a_{2}}^{13/4}\mathrm{AMD}^{2}\left(\sqrt{{a_{2}}}{m_{2}}+\sqrt{{a_{3}}}{m_{3}}\right)^{2}}{256{a_{3}}^{33/4}{m_{0}}^{2}{m_{2}}^{3/2}\sqrt{{m_{3}}}(a+b)}\left(\frac{b}{a}+3\right)+\frac{165{a_{2}}^{9/4}\mathrm{AMD}\left(\sqrt{{a_{2}}}{m_{2}}+\sqrt{{a_{3}}}{m_{3}}\right)}{32{a_{3}}^{19/4}{m_{0}}\sqrt{{m_{2}}}\sqrt{{m_{3}}}}$
@@
${\tilde{D}}_{3}$ $=\frac{675{a_{2}}^{2}\mathrm{AMD}}{32{a_{3}}^{6}{m_{0}}^{2}a}\left(\sqrt{\frac{a_{2}}{a_{3}}}\frac{{m_{3}}}{{m_{2}}}+\frac{{a_{2}}}{2{a_{3}}}+\frac{{m_{3}}^{2}}{2{m_{2}}^{2}}\right)+\frac{3{a_{2}}}{4{a_{3}}^{3}{m_{0}}}\left(\frac{3{a_{2}}{m_{2}}}{{a_{3}}{m_{3}}}-\frac{7{m_{3}}}{4{m_{2}}}\right)\,$,
${\tilde{\Delta}}_{3}$ $=\frac{3{a_{2}}^{3/2}\sqrt{\mathcal{G}}}{8{a_{3}}^{7/2}\sqrt{{m_{0}}}}\left(\sqrt{{a_{2}}}{m_{2}}-\sqrt{{a_{3}}}{m_{3}}\right)-\frac{225{a_{2}}^{2}\mathrm{AMD}^{2}}{32{a_{3}}^{6}{m_{0}}^{2}a}\left(\sqrt{\frac{a_{2}}{a_{3}}}\frac{{m_{3}}}{{m_{2}}}+\frac{{a_{2}}}{2a_{3}}+\frac{{m_{3}}^{2}}{2{m_{2}}^{2}}\right)$
$\phantom{=}+\frac{3{a_{2}}\mathrm{AMD}}{2{a_{3}}^{3}{m_{0}}}\left(\frac{m_{3}}{{m_{2}}}-\frac{{a_{2}}{m_{2}}}{{a_{3}}{m_{3}}}\right)\,$,
where $\tilde{D}(\sigma_{0})=\tilde{D}_{1}\sigma_{0}+\tilde{\Delta}_{1}$ and $\tilde{E}(\sigma_{0})=\tilde{D}_{3}\sigma_{0}+\tilde{\Delta}_{3}$  .
@@
In our numerical example, after dropping the additive constants, as well as an irrelevant term depending only on $\sigma_{0}$, the secular normal form in the octupole approximation resumes the form of Eq. (62) with
${\tilde{A}}\!=0,$ ${\tilde{C}}\!=-0.0283636,$ ${\tilde{B}}\!=-0.029019,$
${\tilde{D}}(\sigma_{0})\!=0.0011798-0.21118\,\sigma_{0},$ ${\tilde{E}}(\sigma_{0})\!=0.00153399-0.270077\,\sigma_{0}\,.$
@@
$\widetilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}=\mathcal{K}_{I}(\tilde{\sigma}_{1},\tilde{\sigma}_{3};\sigma_{0})=\mathcal{A}\tilde{\sigma}_{1}^{2}+\mathcal{C}\tilde{\sigma}_{3}^{2}+(\mathcal{D}_{1}{\sigma}_{0}+\Delta_{1})\tilde{\sigma}_{1}+(\mathcal{D}_{3}{\sigma}_{0}+\Delta_{3})\tilde{\sigma}_{3}$ (63)